# GFlowVLM: Enhancing Multi-step Reasoning in Vision-Language Models with Generative Flow Networks

### v1

@@ Abstract
Vision-Language Models (VLMs) have recently shown promising advancements in sequential decision-making tasks through task-specific fine-tuning. However, common fine-tuning methods, such as Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL) techniques like Proximal Policy Optimization (PPO), present notable limitations: SFT assumes Independent and Identically Distributed (IID) data, while PPO focuses on maximizing cumulative rewards. These limitations often restrict solution diversity and hinder generalization in multi-step reasoning tasks. To address these challenges, we introduce a novel framework, GFlowVLM, a framework that fine-tune VLMs using Generative Flow Networks (GFlowNets) to promote generation of diverse solutions for complex reasoning tasks. GFlowVLM models the environment as a non-Markovian decision process, allowing it to capture long-term dependencies essential for real-world applications. It takes observations and task descriptions as inputs to prompt chain-of-thought (CoT) reasoning which subsequently guides action selection. We use task based rewards to fine-tune VLM with GFlowNets. This approach enables VLMs to outperform prior fine-tuning methods, including SFT and RL. Empirical results demonstrate the effectiveness of GFlowVLM on complex tasks such as card games (NumberLine, BlackJack) and embodied planning tasks (ALFWorld), showing enhanced training efficiency, solution diversity, and stronger generalization capabilities across both in-distribution and out-of-distribution scenarios.
@@ 1 Introduction
Vision-Language Models (VLMs) have achieved remarkable results in generalized tasks such as image captioning and visual question answering [15, 20, 45]. However, they struggle with structured reasoning in sequential decision making tasks that require causal understanding data [3], especially in long horizon planning for tasks such as embodied AI, where agent must capture long term dependencies.
@@
Recent advancements in large language models demonstrate emergent reasoning capabilities by leveraging Chain-of-Thought (CoT) [35] reasoning, that enhances decision-making in multi-step interactive environments [35, 25, 7]. Typically, these models are fine-tuned using specialized visual instruction-following datasets through Supervised Fine Tuning (SFT) methods [20, 18], without active interaction with the environment, or optimized through Reinforcement Learning (RL) approaches, such as Proximal Policy Optimization (PPO) [29, 41]. However, SFT approaches often limits generalization to unseen scenarios, as training relies on maximizing the likelihood over a limited, specialized dataset, thereby restricting diversity in the solution space [17]. Furthermore, for cases where RL methods tend to focus on immediate rewards, can hinder the model’s ability to consider long-term outcomes. Consequently, limited exploration in these models can result in an oversight of more optimal long-term strategies, leading to suboptimal performance in complex tasks [44].
@@
Unlike traditional reinforcement learning (RL) methods, which focus on maximizing cumulative rewards [6, 5], Generative Flow Networks (GFlowNets) [1] train stochastic policies to sample diverse, high-reward sequences (e.g., token sequences) with probabilities proportional to a specified reward function $R(x)$[2]. This approach samples sequences based on the reward function’s distribution, enabling it to find a broader range of high-reward solutions beyond those typically identified by reward-maximizing techniques. Recent studies have applied GFlowNets to multi-step reasoning within the Large Language Models (LLMs) framework, demonstrating their effectiveness over maximum likelihood training and traditional reward-maximization methods [40, 33, 8]. However, these methods lack the multimodal capabilities essential for embodied AI tasks that require integrated visual and textual inputs. Additionally, related works, such as FoR [40], assume Markovian structures within this framework, which may not capture the dependencies necessary for complex, long-horizon, real-world reasoning tasks.
@@
To address these gaps in VLMs, we propose a novel approach GFlowVLM-that leverages GFlowNets for end-to-end fine-tuning in sequential tasks, explicitly modeling non-Markovian flows for richer multimodal reasoning. This framework is, to our knowledge, the first to integrate GFlowNets with VLMs in an end-to-end setting, tackling the unique challenges of multimodal, sequential reasoning essential for complex planning tasks. Our approach initializes policy with a pretrained VLM and fine-tunes it using GFlowNets, guiding VLMs toward structured reasoning processes that capture logical dependencies between successive states. By implicitly representing reasoning as a tree structure—where nodes correspond to states with prior actions and observations, and edges represent actions leading to the next state—GFlowVLM enhances efficient learning of diverse and complex reasoning sequences. Empirical results demonstrate that our method outperforms standard fine-tuning techniques, such as SFT and RL methods like PPO, in achieving structured, multimodal reasoning.
@@ 3 Preliminaries
GFlowNets are models that amortize the cost of sampling from a target distribution over terminal states $\mathcal{X}$ by learning an approximation of this distribution based on its reward function. Given a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{S},\mathcal{A})$ with states $\mathcal{S}$ and directed actions $\mathcal{A}$, there is an initial state $s_{0}$ and terminal states $\mathcal{X}\subset\mathcal{S}$. A trajectory $\tau=(s_{0}\to\ldots\to s_{n})$ represents a complete sequence ending in a terminal state $x\in\mathcal{X}$. The trajectory flow $F:\mathcal{T}\rightarrow\mathbb{R}^{+}$ defines flows over trajectories, with state flow $F(s)=\sum_{s\in\tau}F(\tau)$. A forward policy $P_{F}(\cdot|s)$, often parametrized by a neural network, induces a distribution over trajectories and a marginal distribution over terminal states, with probabilities given by: $P_{F}(\tau)=P_{F}(s_{0}\to\ldots\to s_{n})=\prod_{t=0}^{n-1}P_{F}(s_{t+1}|s_{t})\quad\forall\tau\in\mathcal{T}$. Similarly, a backward policy $P_{B}(\tau)=P_{B}(s_{n}\to\ldots\to s_{0})=\prod_{t=0}^{n-1}P_{B}(s_{t}|s_{t+1})\quad\forall\tau\in\mathcal{T}$. Given a non-negative reward function $R:\mathcal{X}\rightarrow\mathbb{R}^{+}$, GFlowNets aim to estimate a policy where the likelihood of sampling $x\in\mathcal{X}$ is proportional to $R(x)$. Thus, there exists a constant $Z$ such that: $R(x)=Z\sum_{\tau=(s_{0}\rightarrow\ldots\rightarrow s_{n}=x)}P_{F}(\tau)\quad\forall x\in\mathcal{X}$, where $Z=F(s_{0})=\sum_{\tau\in\mathcal{T}}F(\tau)$ is total flow at the initial state.
@@ 3.1 Motivating Experiment
Figure 1: Overview of the prediction of diverse sequence using Gflownets as compared to PPO. The model takes the image of sequence and prompt as input, and generates the next number of sequence by implicitly modeling the causality.
@@
To demonstrate the limitations of traditional approaches and highlight the dependencies captured by GFlowNets, we design a toy experiment combining two types of numerical sequences: $(i)$ the Fibonacci sequence, defined by $F(n)=F(n-1)+F(n-2)$, where each term is the sum of the previous two, and $(ii)$ an arithmetic sequence with a constant increment, $S(n)=S(n-1)+k$, where $k$ is a fixed step size (e.g., $k=2$ for sequences like $[2,4,6,\dots]$). The task presents the model with an image of a partial sequence and a prompt (as shown in Fig. 1), to predict the next number in the sequence. We evaluate the performance of fine-tuning VLM (LLAVA-v1.6-Mistral-7B [19]) using SFT, PPO, and GFlowNets, with temperature parameters $\alpha=1$ and $\alpha>1$ to assess stochastic performance, as shown in Tab. 1. Success rate (SR), measured as the percentage of correct next-number predictions across 1,000 samples, shows GFlowVLM outperform PPO by 26% and generate 40% more diverse solutions. Compared to SFT, GFlowVLM achieves a 54% higher success rate and yield 59% more diversity in responses, underscoring their strength in learning and generalizing causal structures. This advantage stems from GFlowNets’ ability to infer underlying causal reasoning structure of sequence by sequentially sampling reasoning paths, in contrast to the limited diversity observed with SFT and PPO.
@@ 4 Methodology
Figure 2: Overall framework of proposed method: We propose a framework for fine-tuning large VLMs using Gflownets.The input $z_{0:t}$ at time step $t$ consists of a visual observation $o_{t}$ and an input prompt $p_{t}$ containing goal description, history states $s_{0:t}$, history actions $a_{0:t}$, and admissible actions $\mathcal{A}_{t}$, and outputs CoT reasoning $c_{t}$, and action $a_{t}$. The $a_{t}$ is executed in the environment to obtain reward $r_{t}(s_{t},a_{t})$, next observation $o_{t+1}$, and action space $\mathcal{A}_{t+1}$. $f$ generates the next prompt $p_{t+1}$ using description of next observation $o_{t+1}$ (if applicable), history of states $s_{0:t}$ and actions $a_{0:t}$ and next admissible actions $\mathcal{A}_{t+1}$. The sequence of transitions $<s_{t},a_{t},r_{t},c_{t}>$ is added to buffer to update the forward policy $P_{F}$ using Gflownets. $x_{n}$ represents the terminal state of a sequence.
This work utilizes GFlowNet’s structure learning to enhance the VLM’s ability to obtain high-quality, diverse solutions whose distribution is proportional to the reward function. By fine tuning VLMs using GFlowNets, it allows the solutions to be sampled from the distribution of the reward function, which prevents learning policies settled around a small number of modes. Fig. 2 shows the overall pipeline of our proposed framework. The model takes current observation image $o_{t}$ and designed task specific prompt $p_{t}$ as the input. $p_{t}$ contains the description of the goal, history actions $a_{1:t-1}$, history states $s_{1:t-1}$ and admissible action space corresponding to the current observation $o_{t}$. To incorporate Non-Markovian assumption, input $z_{0:t}$ include history actions $a_{0:t}$ and states $s_{0:t}$, respectively along with the input image $o_{t}$. ∗*∗∗*∗*Only the current input image $o_{t}$ is used, as including every intermediate image would be computationally costly, and current VLMs do not perform optimally with multiple images as input.The desired output format includes the CoT reasoning $c_{t}$ and action $a_{t}$, where $a_{t}$ directly interacts with the environment.
@@ 4.1 VLM as a policy: Fine tuning VLMs using GFlowNets to estimate actions
We retain the same prompt format as  [41] for a fair comparison. However, to incorporate historical context in decision-making, we modify the prompt template to include the history of states and actions predicted by the VLM, as shown in Table Tab. 2. The textual prompt $p_{t}$ contains the goal description $g$, the history of states $s_{0:t}$ and actions $a_{0:t}$, and the action space $\mathcal{A}_{t+1}$ available after interacting with the environment. For certain tasks $q$ that may contain observation-dependent information, such as the textual description $d(o_{t+1})$ of the observation $o_{t+1}$, the function $f$ generates the prompt $p_{t+1}$ as: $p_{t+1}=f(d(o_{t+1})\cdot\mathbb{I}_{\{\text{Task}=\text{q}\}},s_{0:t},a_{0:t},\mathcal{A}_{t+1})$, where $\mathbb{I}$ is an indicator function which is 1 only for a certain task $q$ if the observation-dependent information is available.
@@
Before selecting an action at each step $t$, we incorporate a CoT reasoning mechanism, where the model generates intermediate reasoning steps to guide the action selection process. At time $t$, the VLM first generates a reasoning CoT $c_{t}$, which includes a description of the image and intermediate thoughts. Since VLMs are pre-trained on large-scale image-caption data, CoT steps provide additional context and help the model explicitly consider dependencies between different states before selecting the next action. The CoT then guides the action selection. The probabilities for the CoT and action sequences of tokens are defined as follows:
@@
$\sum_{i=1}^{n_{a}}\log P_{\text{VLM}}(w_{i}|w_{<i},c_{t},z_{0:t},g;\theta)$ (4)
where $n_{c}$ and $n_{a}$ represent the number of tokens in the CoT sequence $c_{t}$ and action sequence $a_{t}$, respectively, and $w_{i}$ represents the $i$-th text token in a sequence. Here, $P_{\text{VLM}}(w_{i}|w_{<i},z_{0:t},c_{t},g;\theta)$ and $P_{\text{VLM}}(w_{j}|w_{<j},z_{0:t},g;\theta)$ denote the VLM’s token-level log probabilities for the action and CoT sequences, conditioned on previous tokens, the history of states $z_{0:t}$, the history of actions $a_{0:t-1}$ and goal description $g$. Then, the final forward policy $P_{F}(z_{t+1}|z_{0:t},g;\theta)$ is then computed as a weighted sum of the action probabilities $P_{\text{CoT}}(c_{t}|z_{0:t},g;\theta)$, based on the CoT reasoning, and the original action probabilities $P_{\text{Action}}(a_{t}|z_{0:t},g;\theta)$:
where $\lambda\in[0,1]$ is a weighting factor that controls the influence of the CoT reasoning on the final action selection. The CoT probabilities $P_{\text{CoT}}(c_{t}|z_{0:t},g;\theta)$ provide a structured, intermediate reasoning context that refines the decision-making process, ensuring that the final action is selected with consideration of both direct state information and the model’s internal thought process. We perform an ablation study on the effect of $\lambda$, and we selected $\lambda=0.4$ in our work, as discussed in Supplementary material.

### v2

@@ Abstract
Vision-Language Models (VLMs) have recently shown promising advancements in sequential decision-making tasks through task-specific fine-tuning. However, common fine-tuning methods, such as Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL) techniques like Proximal Policy Optimization (PPO), present notable limitations: SFT assumes Independent and Identically Distributed (IID) data, while PPO focuses on maximizing cumulative rewards. These limitations often restrict solution diversity and hinder generalization in multi-step reasoning tasks. To address these challenges, we introduce a novel framework, GFlowVLM, a framework that fine-tune VLMs using Generative Flow Networks (GFlowNets) to promote generation of diverse solutions for complex reasoning tasks. GFlowVLM models the environment as a non-Markovian decision process, allowing it to capture long-term dependencies essential for real-world applications. It takes observations and task descriptions as inputs to prompt chain-of-thought (CoT) reasoning which subsequently guides action selection. We use task based rewards to fine-tune VLM with GFlowNets. This approach enables VLMs to outperform prior fine-tuning methods, including SFT and RL. Empirical results demonstrate the effectiveness of GFlowVLM on complex tasks such as card games (NumberLine, BlackJack) and embodied planning tasks (ALFWorld), showing enhanced training efficiency, solution diversity, and stronger generalization capabilities across both in-distribution and out-of-distribution scenarios. Project page is available at https://mk322.github.io/gflowvlm/.
@@ 1 Introduction
Vision-Language Models (VLMs) have achieved remarkable results in generalized tasks such as image captioning and visual question answering [15, 20, 45]. However, they struggle with structured reasoning in sequential decision making tasks that require causal understanding [3], especially in long horizon planning for tasks such as embodied AI, where agent must capture long term dependencies.
@@
Recent advancements in LLMs and VLMs demonstrate emergent reasoning capabilities by leveraging Chain-of-Thought (CoT) reasoning, that enhances decision-making in multi-step interactive environments [35, 25, 7]. Typically, these models are fine-tuned using specialized visual instruction-following datasets through Supervised Fine Tuning (SFT) methods [20, 18], without active interaction with the environment, or optimized through Reinforcement Learning (RL) approaches, such as Proximal Policy Optimization (PPO) [29, 41]. However, SFT approaches often limits generalization to unseen scenarios, as training relies on maximizing the likelihood over a limited, specialized dataset, thereby restricting diversity in the solution space [17]. Furthermore, RL methods like PPO tend to prioritize short-term rewards, which can hinder the model’s ability to consider long-term outcomes. Consequently, limited exploration in these models may lead to the oversight of more optimal long-term strategies, resulting in suboptimal performance in complex tasks [44].
@@
In contrast to traditional reinforcement learning (RL) methods, which focus on maximizing cumulative rewards [6, 5], Generative Flow Networks (GFlowNets) [1] train stochastic policies to sample diverse, high-reward sequences (e.g., token sequences) with probabilities proportional to a specified reward function $R(x)$ [2]. This approach samples sequences based on the reward function’s distribution, enabling it to find a broader range of high-reward solutions beyond those typically identified by reward-maximizing techniques. Recent studies have applied GFlowNets to multi-step reasoning within the Large Language Models (LLMs) framework, demonstrating their effectiveness over maximum likelihood training and traditional reward-maximization methods [40, 33, 8]. However, these methods lack the multimodal capabilities which are crucial for embodied AI tasks requiring the integration of visual and textual information. Additionally, related works, such as FoR [40], relies on Markovian structures within this framework, which may fail to capture the long-range dependencies necessary for complex, real-world reasoning tasks.
@@
To address these limitations, we introduce GFlowVLM, a novel approach integrating GFlowNets with VLMs in an end-to-end fine-tuning framework. It explicitly models non-Markovian flows, enabling richer multimodal reasoning suitable for complex sequential decision-making. To our knowledge, GFlowVLM is the first to fuse GFlowNets with VLMs directly, addressing the distinct challenges posed by multimodal, sequential reasoning crucial in structured planning environments. Our approach initializes a policy with a pretrained VLM and fine-tunes it using GFlowNets, guiding VLMs toward structured reasoning processes that capture logical dependencies between successive states. By implicitly representing reasoning as a tree structure—where nodes correspond to states with prior actions and observations, and edges represent actions leading to the next state—GFlowVLM enhances efficient learning of diverse and complex reasoning sequences. Empirical results demonstrate that GFlowVLM outperforms standard fine-tuning techniques including SFT and RL methods including PPO, by enhancing structured multimodal reasoning capabilities.
@@ 3 Preliminaries
GFlowNets are models that amortize the cost of sampling from a target distribution over terminal states $\mathcal{X}$ by learning an approximation of this distribution based on its reward function. Given a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{S},\mathcal{A})$ with states $\mathcal{S}$ and directed actions $\mathcal{A}$, there is an initial state $s_{0}$ and terminal states $\mathcal{X}\subset\mathcal{S}$. A trajectory $\tau=(s_{0}\to\ldots\to s_{n}=x)$ represents a complete sequence ending in a terminal state $x\in\mathcal{X}$. The trajectory flow $F:\mathcal{T}\rightarrow\mathbb{R}_{\geq 0}$ defines flows over trajectories, with state flow $F(s)=\sum_{s\in\tau}F(\tau)$. A forward policy $P_{F}(\cdot|s)$, often parametrized by a neural network, induces a distribution over trajectories and a marginal distribution over terminal states, with probabilities given by: $P_{F}(\tau)=P_{F}(s_{0}\to\ldots\to s_{n})=\prod_{t=0}^{n-1}P_{F}(s_{t+1}|s_{t})\quad\forall\tau\in\mathcal{T}$. Similarly, a backward policy $P_{B}(\tau)=P_{B}(s_{n}\to\ldots\to s_{0})=\prod_{t=0}^{n-1}P_{B}(s_{t}|s_{t+1})\quad\forall\tau\in\mathcal{T}$. Given a non-negative reward function $R:\mathcal{X}\rightarrow\mathbb{R}_{\geq 0}$, GFlowNets aim to estimate a policy where the likelihood of sampling $x\in\mathcal{X}$ is proportional to $R(x)$. Thus, there exists a constant $Z$ such that: $R(x)=Z\sum_{\tau=(s_{0}\rightarrow\ldots\rightarrow s_{n}=x)}P_{F}(\tau)\quad\forall x\in\mathcal{X}$, where $Z=F(s_{0})=\sum_{\tau\in\mathcal{T}}F(\tau)$ is total flow at the initial state. See Appendix A for more details.
@@ 3.1 Motivating Experiment
Figure 1: Overview of the prediction of diverse sequence using Gflownets as compared to PPO. The model takes the image of sequence and prompt as input, and generates the next number of sequence by implicitly modeling the causality. See Fig. 4 for a practical example.
@@
To demonstrate the limitations of traditional approaches and highlight the dependencies captured by GFlowNets, we design a toy experiment combining two types of numerical sequences: $(i)$ the Fibonacci sequence, defined by $F(n)=F(n-1)+F(n-2)$, where each term is the sum of the previous two, and $(ii)$ an arithmetic sequence with a constant increment, $S(n)=S(n-1)+k$, where $k$ is a fixed step size (e.g., $k=2$ for sequences like $[2,4,6,\dots]$). The task presents the model with an image of a partial sequence and a prompt (as shown in Fig. 1), to predict the next number in the sequence. We evaluate the performance of fine-tuning VLM (LLAVA-v1.6-Mistral-7B [19]) using SFT, PPO, and GFlowNets, with temperature parameters $\alpha=1$ and $\alpha>1$ to assess stochastic performance, as shown in  Tab. 1. Success rate (SR), measured as the percentage of correct next-number predictions across 1,000 samples, shows GFlowVLM outperform PPO by 26% and generate 40% more diverse solutions. Compared to SFT, GFlowVLM achieves a 54% higher success rate and yield 59% more diversity in responses, underscoring their strength in learning and generalizing causal structures. This advantage stems from GFlowNets’ ability to infer underlying causal reasoning structure of sequence by sequentially sampling reasoning paths, in contrast to the limited diversity observed with SFT and PPO. While this toy example highlights key conceptual benefits of GFlowNets, we include a practical example in Sec. A.2 of Supplementary Material demonstrating how GFlowVLM can be applied to embodied AI tasks in ALFWorld, showcasing its real-world reasoning capabilities. This addition provides further evidence of the method’s utility beyond synthetic settings and illustrates its effectiveness in a more grounded, task-oriented scenario.
@@ 4 Methodology
Figure 2: Overall framework of proposed method: The input $z_{0:t}$ at time step $t$ consists of a visual observation $o_{t}$ and an input prompt $p_{t}$ containing goal description, history states $s_{0:t}$, history actions $a_{0:t}$, and admissible actions $\mathcal{A}_{t}$, and outputs CoT reasoning $c_{t}$, and action $a_{t}$. The $a_{t}$ is executed in the environment to obtain reward $r_{t}(s_{t},a_{t})$, next observation $o_{t+1}$, and action space $\mathcal{A}_{t+1}$. $f$ generates the next prompt $p_{t+1}$ using description of next observation $o_{t+1}$ (if applicable), history of states $s_{0:t}$ and actions $a_{0:t}$ and next admissible actions $\mathcal{A}_{t+1}$. This generates multiple trajectories. The transitions $<s_{t},a_{t},r_{t},c_{t}>$ , $<s_{t}^{\prime},a_{t}^{\prime},r_{t}^{\prime},c_{t}^{\prime}>$ and $<s_{t}^{\prime\prime},a_{t}^{\prime\prime},r_{t}^{\prime\prime},c_{t}^{\prime\prime}>$ across different trajectories are added to buffer to update the forward policy $P_{F}$ using GFlowNets. $\{x,x^{\prime},x^{\prime\prime}\}\in\mathcal{X}$ represent the terminal states of sequences. $R(x)$ represents the non-negative reward obtained from the environment (after reward shaping, if applicable) at terminal state $x$ of a trajectory.
This work utilizes GFlowNet’s structure learning to enhance the VLM’s ability to obtain high-quality, diverse solutions whose distribution is proportional to the reward function. By fine tuning VLMs using GFlowNets, it allows the solutions to be sampled from the distribution of the reward function, which prevents learning policies settled around a small number of modes. Fig. 2 shows the overall pipeline of our proposed framework. The model takes current observation image $o_{t}$ and designed task specific prompt $p_{t}$ as the input. $p_{t}$ contains the description of the goal, history actions $a_{1:t-1}$, history states $s_{1:t-1}$ and admissible action space corresponding to the current observation $o_{t}$. To incorporate non-Markovian assumption, input $z_{0:t}$ include history actions $a_{0:t}$ and states $s_{0:t}$, respectively along with the input image $o_{t}$∗*∗∗*∗*Only the current input image $o_{t}$ is used, as including every intermediate image would be computationally costly, and current VLMs do not perform optimally with multiple images as input.. The desired output format includes the CoT reasoning $c_{t}$ and action $a_{t}$, where $a_{t}$ directly interacts with the environment.
@@ 4.1 VLM as a policy: Fine tuning VLMs using GFlowNets to estimate actions
We retain the same prompt format as  [41] for a fair comparison. However, to incorporate historical context in decision-making, we modify the prompt template to include the history of states and actions predicted by the VLM, as shown in Tab. 2. The textual prompt $p_{t}$ contains the goal description $g$, the history of states $s_{0:t}$ and actions $a_{0:t}$, and the action space $\mathcal{A}_{t+1}$ available after interacting with the environment. For certain tasks $q$ that may contain observation-dependent information, such as the textual description $d(o_{t+1})$ of the observation $o_{t+1}$, the function $f$ generates the prompt $p_{t+1}$ as: $p_{t+1}=f(d(o_{t+1})\cdot\mathbb{I}_{\{\text{Task}=\text{q}\}},s_{0:t},a_{0:t},\mathcal{A}_{t+1})$, where $\mathbb{I}$ is an indicator function which is 1 only for a certain task $q$ if the observation-dependent information is available.
@@
Before selecting an action at each step $t$, we incorporate a CoT reasoning mechanism, where the model generates intermediate reasoning steps to guide the action selection process. At time $t$, the VLM first generates a reasoning CoT $c_{t}$, which includes a description of the image and intermediate thoughts. Since VLMs are pre-trained on large-scale image-caption data, CoT steps provide additional context and help the model explicitly consider dependencies between different states before selecting the next action. The CoT then guides the action selection. We define $z_{t}$ as the structured state at time $t$ in the trajectory, which includes the action $a_{t}$, environment state $s_{t}$, and visual observation $o_{t}$. More precisely, the trajectory $z_{0:t}$ is composed of the current visual observation $o_{t}$ and an input prompt $p_{t}$ that contains the goal description $g$, the history of environment states $s_{0:t-1}$, actions $a_{0:t-1}$, and the set of admissible actions $\mathcal{A}_{t}$. See Sec. C.1 for details.
The probabilities for the CoT and action sequences of tokens are defined as follows:
@@
$\prod_{i=1}^{n_{a}}P_{\text{VLM}}(w_{i}|w_{<i},c_{t},z_{0:t},g;\theta)$ (4)
where $n_{c}$ and $n_{a}$ represent the number of tokens in the CoT sequence $c_{t}$ and action sequence $a_{t}$, respectively, and $w_{i}$ represents the $i$-th text token in a sequence. Here, $P_{\text{VLM}}(w_{i}|w_{<i},c_{t},z_{0:t},g;\theta)$ and $P_{\text{VLM}}(w_{j}|w_{<j},z_{0:t},g;\theta)$ denote the VLM’s token-level likelihoods for the action and CoT sequences, conditioned on previous tokens, the history of states $z_{0:t}$, and goal description $g$. The log forward policy $\log P_{F}(z_{t+1}|z_{0:t},g;\theta)$ is then computed as a weighted sum of the log probabilities of CoT tokens $\log P_{\text{CoT}}(c_{t}|z_{0:t},g;\theta)$, and the original log action probabilities $\log P_{\text{Action}}(a_{t}|z_{0:t},g;\theta)$:
$\begin{split}\log P_{F}(z_{t+1}|z_{0:t},g;\theta)=\log P_{\text{Action}}(a_{t}|z_{0:t},c_{t},g;\theta)+\\ \lambda\log P_{\text{CoT}}(c_{t}|z_{0:t},g;\theta),\end{split}$ (5)
where $\lambda\in[0,1]$ is a weighting factor that controls the influence of the CoT reasoning on the final action selection. The CoT probabilities $P_{\text{CoT}}(c_{t}|z_{0:t},g;\theta)$ provide a structured, intermediate reasoning context that refines the decision-making process, ensuring that the final action is selected with consideration of both direct state information and the model’s internal thought process. We perform an ablation study on the effect of $\lambda$, and we selected $\lambda=0.4$ in our work, as discussed in the Sec. D.5 of Supplementary Material.